OpenStudy (anonymous):

Find exact value given that secx = 3/2, cscy = 3, and x and y are in Quadrant I. sin(x+y)

OpenStudy (anonymous):

So far I have \[\frac{ 2\sqrt{10} }{ 3 }+\frac{ 2 }{ 9 }\]

OpenStudy (anonymous):

you got a bunch of work to do

OpenStudy (anonymous):

\[\sin(x+y)=\sin(x)\cos(y)+\cos(x)\sin(y)\]

OpenStudy (anonymous):

at the moment you know \[\cos(x)=\frac{2}{3}\] and \[\sin(y)=\frac{1}{3}\]

OpenStudy (anonymous):

you need two more numbers, \[\sin(x)\] and \[\cos(y)\]

OpenStudy (anonymous):

sin x = sqrt(5)/3 and cos y = 2sqrt(2)/3

OpenStudy (anonymous):

ok then you are in good shape plug them in and you are done

OpenStudy (anonymous):

you know all four numbers for \[\sin(x)\cos(y)+\cos(x)\sin(y)\]

OpenStudy (anonymous):

I did. But I don't know where to go from 2sqrt10/3 + 2/9

OpenStudy (anonymous):

lets see

OpenStudy (anonymous):

\[\frac{\sqrt5}{3}\times \frac{2\sqrt2}{3}+\frac{2}{3}\times \frac{1}{3}\] \[=\frac{2\sqrt{10}+2}{9}\] if my arithmetic is correct

OpenStudy (anonymous):

Yes, you're right. I did mine wrong.

OpenStudy (anonymous):

i think you made a mistake in the denominator of the first part

OpenStudy (anonymous):

other than that it is correct

OpenStudy (anonymous):

Awesome, thanks

OpenStudy (anonymous):

yw